OpenStudy (anonymous):

If sinxsiny=1/4 and 3tanx=4tany , then prove that sin(x+y)=7/16 (wrong question)

OpenStudy (anonymous):

sinxsiny=1/4 ..i 3tanx=4tany ....ii =4tany=3tanx ...iii Dividing i by ii, cosx siny/3 = cosy/16siny siny cosx =3cosy /16siny ....iv Dividing i by iii, cosysinx/4 = cosx/12sinx sinx cosy =cosx /3 siny.....v

OpenStudy (anonymous):

sin(x+y) = sinx cosy + sinycosx =cosx/3sinx + 3cosy/16siny =1/3tanx + 3/16tany =1/3tanx + 3/4(3tanx) =1/3tanx + 1/4tanx =7/12tanx

OpenStudy (anonymous):

sin(x+y)=7/16tany

OpenStudy (phi):

This is not going to happen If sinxsiny=1/4 and 3tanx=4tany , then prove that sin(x+y)=7/16 you can solve numerically to find sin(x+y) = 0.870460525 which is close to 0.875 = 7/8

OpenStudy (phi):

you can use trig substitutions and algebra to find angles x and y